OpenStudy (anonymous):

You are traveling to your mother’s house that is 234 miles away. If you are currently twice as far from your home as you are from your mother’s house, how far have you traveled? Round your answer to the nearest tenth mile

OpenStudy (anonymous):

Let distance from mother house =x then, distance traveled by me (i.e distance from my home)= 2x x + 2x = 234 [According to question] 3x= 234 => x = 234/3= 78 distance traveled by me = 2* 78 = 156 miles

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

let x - distance from your house then 2x would be the distance from your mother's house (since you are twice as far from your mother's house) The total distance is 234 miles which is the sum of the distance from your house and the distance from your mothers house. So: x + 2x = 234 3x = 234 x = 234/3